OpenStudy (anonymous):

Solve x/4 + 12 < 42. x < 7.5 x < 120 x < 13.5 x < 216 Solve x/4 + 12 < 42. x < 7.5 x < 120 x < 13.5 x < 216 @Mathematics

OpenStudy (across):

Solve it exactly the same way you would a normal equation, that is,\[\frac{x}{4}+12=42.\]

OpenStudy (anonymous):

which is

OpenStudy (anonymous):

x/4=42-12 and keep going :-)

OpenStudy (anonymous):

so 120

OpenStudy (anonymous):

Yes. You always can check answer by plugging it into original equation.